Limestone (πΆπ‘ŽπΆπ‘‚3) dissolves in hydrochloric acid (𝐻𝐢𝑙) for form calcium chloride (πΆπ‘ŽπΆπ‘™2), carbon dioxide, and water. How many moles of HCl are required to dissolve 205g of limestone? How many moles of water are formed?

Show solution step by step with explaination



1
Expert's answer
2020-08-31T12:46:42-0400

Write the reaction:

CaCO3+2HCl→CaCl2+CO2+H2O.

We have 205 g of chemically pure limestone. How many moles is this?


"n_l=\frac{m}{\Mu}=\frac{205}{40+12+16\cdot3}=2.05\text{ mol}."


What is a mole? It's the amount of substance that contains a specific number of particles. Therefore, 2.05 moles of any substance contains equal number of particles. This is the key to proceed: we need exactly 2.05 moles of HCl to react with limestone. But look at the reaction: it says that one molecule of CaCO3 requires two molecules of HCl, therefore, we need two times more particles for HCl, or 4.1 moles. This is equivalent to the mass of acid


"m=n\Mu=4.1\cdot(1+35.5)=149.7\text{ g}."

According to the reaction, one molecule of limestone produces one molecule of water, thus, for water


"n=2.05\text{ mol},\\
m=n\Mu=2.05(1\cdot2+16)=36.9\text{ g}."